OpenStudy (anonymous):

in an explicit formula an=a1+(n-1)*d , what does the "d" stand for?

OpenStudy (princeharryyy):

common difference between the successive terms

OpenStudy (princeharryyy):

@maggieabc32 :p

OpenStudy (anonymous):

so what would be the successive terms?

OpenStudy (princeharryyy):

suppose thers an ap like 4, 8, 12, 16, ...... and so on..... d between each of the two terms like => 8-4=4; 12-8=4; 16-12 =4; and so on which is same :)

OpenStudy (anonymous):

oh ok i get it! thanks!

OpenStudy (princeharryyy):

:p